OpenStudy (omgthatgirl):

what is I 12-2i I last question please help

OpenStudy (omgthatgirl):

one more wrong answer and i make a 70 :/

OpenStudy (rizags):

do you want the answer or an explanation or both

OpenStudy (omgthatgirl):

both please

OpenStudy (omgthatgirl):

@Rizags

OpenStudy (omgthatgirl):

?

OpenStudy (rizags):

ok, so the magnitude of a vector in the imaginary plane, given coordindates a+bi, is \[\sqrt{a^{2}+b ^{2}}\]

OpenStudy (rizags):

try solving from here

OpenStudy (omgthatgirl):

144+4i?

OpenStudy (omgthatgirl):

@Rizags

OpenStudy (omgthatgirl):

Sorry can you make it fast cause it's late and im really tired from work and want to get some rest :)

OpenStudy (omgthatgirl):

@Rizags

OpenStudy (rizags):

ok sorry, um yea plug in 12 for a and 2 for b, to get \[\sqrt{144+4}=\sqrt{138}=2\sqrt{37}\]

OpenStudy (rizags):

medal fan plz!